OpenStudy (anonymous):

The perimeter of a rectangle is 44 inches. If the width of the rectangle is 7 inches, what is the length? 37 inches 15 inches 22 inches 30 inches

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

 7+7+2x=44 14+2x=44 2x+30 x=15 The length of the rectangle is 15

OpenStudy (anonymous):

So the answer is B.15

OpenStudy (anonymous):

You are awesome!